# tablenotes

### v1

@@ Abstract
Neural Network Potentials (NNPs) have attracted significant attention as a method for accelerating density functional theory (DFT) calculations. However, conventional NNP models typically do not incorporate spin degrees of freedom, limiting their applicability to systems where spin states critically influence material properties, such as transition metal oxides. This study introduces SpinMultiNet, a novel NNP model that integrates spin degrees of freedom through multi-task learning. SpinMultiNet achieves accurate predictions without relying on correct spin values obtained from DFT calculations. Instead, it utilizes initial spin estimates as input and leverages multi-task learning to optimize the spin latent representation while maintaining both $E(3)$ and time-reversal equivariance. Validation on a dataset of transition metal oxides demonstrates the high predictive accuracy of SpinMultiNet. The model successfully reproduces the energy ordering of stable spin configurations originating from superexchange interactions and accurately captures the rhombohedral distortion of the rocksalt structure. These results pave the way for new possibilities in materials simulations that consider spin degrees of freedom, promising future applications in large-scale simulations of various material systems, including magnetic materials.
@@ 1 Introduction
However, most conventional NNPs do not account for spin degrees of freedom, limiting their application to material systems where spin states play a critical role in determining properties, such as transition metal oxides (TMOs). TMOs are known to exhibit diverse magnetic properties due to the presence of transition metal ions with partially filled d-orbitals, and incorporating spin degrees of freedom is crucial for understanding their properties [14]. For example, accurate prediction of the energy difference between ferromagnetic (FM) and antiferromagnetic (AFM) states requires proper representation of the potential energy surface depending on the spin configuration. Recently, SpinGNN [15] and xDeepH [16] have been proposed as NNP models incorporating spin degrees of freedom. These models take spin values as input in addition to atomic configurations and predict spin-dependent potential energies. However, these models require correct spin inputs during prediction, limiting their applicability since correct spin values are often unavailable in realistic scenarios.
To overcome this limitation, this study presents SpinMultiNet, a novel NNP model which utilizes initial spin estimates as input and accurately predicts the spin-dependent potential energy surface. Our model employs multi-task learning to predict energy, forces, and spin simultaneously. This allows the spin latent features to be optimized within the network, enabling highly accurate predictions, even if the input spin is an initial estimate provided by the user. Furthermore, our model is designed to satisfy not only $E(3)$ equivariance but also time-reversal equivariance, which ensures consistent and physically meaningful predictions. These equivariance contribute to improved data efficiency and enhanced generalization capabilities. The main contributions of this work are as follows:
@@ 2 Related Work
In recent years, several NNP models that take spin degrees of freedom into account have been proposed. Magnetic moment tensor potential [17] introduces spin degrees of freedom into the moment tensor potential [18], enabling the learning of spin-dependent potentials. Similarly, mHDNNP [19] proposes a model that incorporates spin interactions into atom-centered symmetry functions. However, these methods are limited to collinear spins. On the other hand, SpinGNN [15] addresses the noncollinear spin by using the inner product of two noncollinear spins as input to the GNN. SpinGNN leverages the high expressive power of GNNs to construct accurate potential models.
A critical aspect of spin-dependent NNP models is ensuring time-reversal equivariance. Time-reversal equivariance describes how the state of a system changes under the time-reversal operation and is essential for the physically accurate handling of latent spin features. For example, under the time-reversal operation, spin flips its sign while energy remains invariant. SpinGNN ensures the time-reversal invariance of the energy by restricting spin-derived features to scalars only. On the other hand, xDeepH [16] achieves physical consistency and higher representational capacity by designing an architecture that is equivariant to the time-reversal operation.
However, existing spin-dependent NNP models still face the challenge of requiring correct spin values as input during prediction. Although DFT calculations can provide correct spin values, relying on them for every NNP prediction is impractical because of the significant computational costs. Alternatively, NNP models capable of calculating the gradient of energy with respect to spin, i.e., magnetic forces [20], can be used to obtain correct spin values in the ground state by optimizing the spin configuration to minimize the energy. However, this optimization process is computationally expensive as it requires repeated calculations using optimization algorithms, making it a significant challenge.
@@
SpinMultiNet accurately predicts energies and spin values even from initial spin estimates, without relying on correct spin values obtained from DFT or spin optimization calculations. This is achieved by performing multi-task learning that simultaneously predicts energy and spin while optimizing spin features in a time-reversal equivariant manner. This approach enables efficient calculation of energies for various spin configurations, addressing the computational cost challenges of conventional methods. Table 1 summarizes the relationship between spin input and output for existing research and SpinMultiNet.
Table 1: Comparison of spin input and output between existing research and SpinMultiNet.
Spin input Spin output
SpinGNN Correct values a
CHGNet Correct values (only one ordering)
Ours (SpinMultiNet) Initial estimates Correct values
  • a
    Obtained from DFT calculations, etc.
@@ 5 Conclusion
In this study, we developed SpinMultiNet, a novel multitasking NNP model which explicitly incorporates spin degrees of freedom. This model can simultaneously predict accurate energies and spin values using initial spin estimates as input, without relying on correct spin values obtained from DFT calculations. This was achieved by employing multi-task learning to simultaneously predict energy and spin, optimizing the latent representation of spin in the process. SpinMultiNet accurately captures the spin-dependent energy landscape and can reproduce important physical phenomena such as superexchange interactions. This paves the way for large-scale simulations of various material systems, including magnetic materials, which were challenging for conventional NNP models.

### v2

@@ Abstract
Neural Network Potentials (NNPs) have attracted significant attention as a method for accelerating density functional theory (DFT) calculations. However, conventional NNP models typically do not incorporate spin degrees of freedom, limiting their applicability to systems where spin states critically influence material properties, such as transition metal oxides. This study introduces SpinMultiNet, a NNP model that integrates spin degrees of freedom through multi-task learning. SpinMultiNet achieves accurate predictions without relying on correct spin values obtained from DFT or spin optimization calculations. Instead, it utilizes initial spin estimates as input and leverages multi-task learning to optimize the spin latent representation while maintaining both $E(3)$ and time-reversal equivariance. Validation on a dataset of transition metal oxides demonstrates the high predictive accuracy of SpinMultiNet. The model successfully reproduces the energy ordering of stable spin configurations originating from superexchange interactions and accurately captures the rhombohedral distortion of the rocksalt structure. These results pave the way for new possibilities in materials simulations that consider spin degrees of freedom, promising future applications in large-scale simulations of various material systems, including magnetic materials.
@@ 1 Introduction
However, most conventional NNPs do not account for spin degrees of freedom, limiting their application to material systems where spin states play a critical role in determining properties, such as transition metal oxides (TMOs). TMOs are known to exhibit diverse magnetic properties due to the presence of transition metal ions with partially filled d-orbitals, and incorporating spin degrees of freedom is crucial for understanding their properties [14]. For example, accurate prediction of the energy difference between ferromagnetic (FM) and antiferromagnetic (AFM) states requires proper representation of the potential energy surface depending on the spin configuration. Recently, several NNP models incorporating spin degrees of freedom have been proposed [15, 16, 17]. These models can predict spin-dependent potential energy by using not only atomic coordinates but also spin values as inputs. However, to fully exploit the potential of these architectures, correct spin information is required as input. In practical situations, obtaining such correct spin values is often challenging, which hinders the utilization of these models.
To overcome this limitation, this study presents SpinMultiNet, a NNP model which utilizes initial spin estimates as input and accurately predicts the spin-dependent potential energy surface. Our model employs multi-task learning to predict energy, forces, and spin simultaneously. This allows the spin latent features to be optimized within the network, enabling highly accurate predictions, even if the input spin is an initial estimate provided by the user. Furthermore, our model is designed to satisfy not only $E(3)$ equivariance but also time-reversal equivariance, which ensures consistent and physically meaningful predictions. These equivariance contribute to improved data efficiency and enhanced generalization capabilities. The main contributions of this work are as follows:
@@ 2 Related Work
In recent years, several NNP models that take spin degrees of freedom into account have been proposed. Magnetic moment tensor potential [15] introduces spin degrees of freedom into the moment tensor potential [18], enabling the learning of spin-dependent potentials. Similarly, mHDNNP [16] proposes a model that incorporates spin interactions into atom-centered symmetry functions. However, these methods are limited to collinear spins. On the other hand, SpinGNN [19] addresses noncollinear spins by extracting scalar features from two noncollinear spin vectors and using these features as input to the GNN. SpinGNN leverages the high expressive power of GNNs to construct accurate potential models. These models have achieved success in several material systems, demonstrating the capabilities of spin-dependent NNPs.
A critical aspect of spin-dependent NNP models is ensuring time-reversal equivariance. Time-reversal equivariance describes how the state of a system changes under the time-reversal operation and is essential for the physically accurate handling of latent spin features. For example, under the time-reversal operation, spin flips its sign while energy remains invariant. SpinGNN ensures the time-reversal invariance of the energy by restricting spin-derived features to time-reversal scalars only. On the other hand, xDeepH [17] achieves physical consistency and higher representational capacity by designing an architecture that is equivariant to the time-reversal operation.
However, most spin-dependent NNP models face the challenge of requiring correct spin values as input during prediction in order to maximize the performance of the architecture. Correct spin values can be obtained through methods such as DFT calculations, iterative energy optimization using magnetic forces [20], or by employing another machine learning model. However, executing these approaches for each NNP prediction is computationally expensive.
@@
SpinMultiNet accurately predicts energies and spin values even from initial spin estimates, without relying on correct spin values obtained from DFT or spin optimization calculations. This is achieved by performing multi-task learning that simultaneously predicts energy and spin while optimizing spin features in a time-reversal equivariant manner. This approach enables efficient calculation of energies for various spin configurations, addressing the computational cost challenges of conventional methods. It should be mentioned that our approach is very similar to the recently proposed SpinGNN++[22]. However, this study specifically focuses on investigating the influence of spin prediction from initial estimates as a supervisory signal on the overall performance of the model.
@@ 5 Conclusion
In this study, we developed SpinMultiNet, a multitasking NNP model which explicitly incorporates spin degrees of freedom. This model can simultaneously predict accurate energies and spin values using initial spin estimates as input, without relying on correct spin values obtained from DFT calculations. This was achieved by employing multi-task learning to simultaneously predict energy and spin, optimizing the latent representation of spin in the process. SpinMultiNet accurately captures the spin-dependent energy landscape and can reproduce important physical phenomena such as superexchange interactions. This paves the way for large-scale simulations of various material systems, including magnetic materials, which were challenging for conventional NNP models.